120 psychology students took a standardized test. The scores are summarized in the GFDT below.

      Scores          Frequency   
140 - 159 10
160 - 179 17
180 - 199 10
200 - 219 18
220 - 239 13
240 - 259 52



The scores are also described in the cumulative table shown below.

      Scores          Frequency   
less than 160 10
less than 180 27
less than 200 37
less than 220
less than 240 68
less than 260 120



What is the missing value in the cumulative GFDT?
Answer #1

missing value in the cumulative GFDT =cumulative frequency till less then 200+frequency in range 200-219 =37+18=55